OpenStudy (anonymous):

Factor: x^3+15x^2+75x+125

OpenStudy (anonymous):

Use factor theorem.

OpenStudy (campbell_st):

well the equation seems to fit the binomial expansion of a cubic \[(a + b)^3 = a^3 + 3a^2b + 3b^2a + b^3\] you just need to find a and b

OpenStudy (anonymous):

no its not difference/sum of cubes

OpenStudy (campbell_st):

no its not the sum of 2 cubes if it was the form would be \[a^3 + b^3 = (a + b)(a^2 - ab + b^2)\]

OpenStudy (anonymous):

You could use the rational roots theorem and the descartes method, then use synthetic division.

OpenStudy (campbell_st):

the the of 2 cubes contains 2 terms... the expression above contains 4.

OpenStudy (campbell_st):

why make it harder than it is... just compare it to the binomial expansion.